If 3.65g of butane is burned underneath a cup holding 1.00 L of water at 21.0 C, what will the final temperature of the water be (ΔHcomb = -3325 kJ/mol)?
1
Expert's answer
2020-11-18T13:45:04-0500

"C_4H_{10} + \frac{13}{2}O_2 \to 4CO_2 + 5H_2O"


3.65g of Butane = 3.56/58 = 0.063 mol

Heat evolved = n(∆"H_c") = 0.063(-3325) = -209.5 kJ


Q = mc∆T = mc("T_2 -T_1")

V = 1.0 L

m = density × volume = 1kg/L × 1.0 L

= 1kg

Q = 209.5 kJ = 209,500J

"T_1" = 21.0°C

"T_2" = T

c = 4200J/kg.°C


209,500 = 1 × 4200 × (T - 21)

209500/4200 = T - 21

49.9 = T - 21

T = 70.9°C